OpenStudy (anonymous):

ln\(e\) = ln\( sqrt(2)/x \) -ln\(e\)

OpenStudy (anonymous):

\[\ln(e)=1\]

OpenStudy (anonymous):

you get \[\ln(\frac{\sqrt{2}}{x})=2\]